# Asymptotic expansions for Legendre functions via differential equations having coalescing turning points

### v2

@@ 3 LG expansions for parabolic cylinder functions
Here we present relevant results for parabolic cylinder functions, and derive LG expansions for the ones that we shall use in our uniform asymptotic expansions for solutions of (1).
Firstly, the following Wronskians hold (see [3, Eqs. 12.2.11 and 12.2.12])
@@
Connection formulas are given as follows ([3, Eqs. 12.2.13, 12.2.18, 12.2.19 and 12.2.20])
@@
Thus, for the coefficients [4, Eq. (1.9)], let us denote $E_{s}(\xi)=e_{s}(\hat{\alpha},\hat{\beta})$. Then we find, using Eqs. 51, 52, 54, and 55, and [4, Eqs. (1.10) - (1.12)], that
@@
(59) $\hat{W}_{n,j}(u,\hat{\alpha},\zeta)=\exp\left\{u\hat{\xi}+\sum\limits_{s=1}^{n-1}{\frac{e_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\hat{\eta}_{n,j}(u,\hat{\alpha},\zeta)\right\}\quad(j=1,3),$
(60) $\hat{W}_{n,2}(u,\hat{\alpha},\zeta)=\exp\left\{-u\hat{\xi}+\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{e_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\hat{\eta}_{n,2}(u,\hat{\alpha},\zeta)\right\},$
@@
(61) $\hat{W}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)=\exp\left\{-u\hat{\xi}+\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{e_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\hat{\eta}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)\right\}.$
Similarly to (28), the LG solutions $\hat{W}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)$ only differ by a multiplicative constant, and their $\zeta$ regions of validity $\Theta_{4}^{\pm}$ (where $\hat{\eta}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)=\mathcal{O}(u^{-n})$ as $u\to\infty$) differ.
@@
Matching solutions recessive at $\zeta=\pm\infty$ and $\zeta=\pm i\infty$, and using Eqs. 15 and 50, provides the following:
(65) $U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)=\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\frac{e^{\frac{1}{4}(u\hat{\alpha}^{2}+1)\pi i}}{\left\{2u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\hat{W}_{n,1}(u,\hat{\alpha},\zeta),$
(66) $U\left(\tfrac{1}{2}u\hat{\alpha}^{2},i\sqrt{2u}\,\zeta\right)=\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\frac{e^{-\frac{1}{4}(u\hat{\alpha}^{2}+1)\pi i}}{\left\{2u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\hat{W}_{n,3}(u,\hat{\alpha},\zeta),$
(67) $U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)=\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\frac{1}{\left\{2u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\hat{W}_{n,2}(u,\hat{\alpha},\zeta),$
@@
(68) $U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta\right)=\pm ie^{\mp\frac{1}{2}u\hat{\alpha}^{2}\pi i}\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\\ \times\frac{1}{\left\{2u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\hat{W}_{n,4}^{\pm}(u,\hat{\alpha},\zeta),$
@@
Next, for derivatives from (13), we arrive at
(69) $\frac{d^{2}\tilde{w}}{d\zeta^{2}}=\left\{u^{2}\left(\zeta^{2}-\hat{\alpha}^{2}\right)+\frac{2\zeta^{2}+\hat{\alpha}^{2}}{\left(\zeta^{2}-\hat{\alpha}^{2}\right)^{2}}\right\}\tilde{w},$
whose solutions are given by
(70) $\tilde{w}=(\zeta^{2}-\hat{\alpha}^{2})^{-1/2}\partial w/\partial\zeta,$
where $w=U(-\frac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta)$ or any other solution of (13). Now, again with $f=\zeta^{2}-\hat{\alpha}^{2}$, but instead of $g=0$ now with $g=(2\zeta^{2}+\hat{\alpha}^{2})/(\zeta^{2}-\hat{\alpha}^{2})^{2}$, we have in place of (51)
@@
Then, similarly to Eqs. 56, 57, and 58, we obtain
@@
with $\hat{\beta}$ also given by (52) and subsequent coefficients by (58). Again, $\tilde{e}_{2s}(\hat{\alpha},\hat{\beta})$ are meromorphic at $\zeta=\pm\hat{\alpha}$ when regarded as functions of $\zeta$, and $\tilde{e}_{2s+1}(\hat{\alpha},\hat{\beta})$ when regarded as functions of $\zeta$ are analytic at all points away from the interval $[-\hat{\alpha},\hat{\alpha}]$.
The LG solutions to (69) are given by $\tilde{w}_{n,j}(u,\hat{\alpha},\zeta)=(\zeta^{2}-\hat{\alpha}^{2})^{-1/4}\tilde{W}_{n,j}(u,\hat{\alpha},\zeta)$ where
(74) $\tilde{W}_{n,j}(u,\hat{\alpha},\zeta)=\exp\left\{u\hat{\xi}+\sum\limits_{s=1}^{n-1}{\frac{\tilde{e}_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\tilde{\eta}_{n,j}(u,\hat{\alpha},\zeta)\right\}\quad(j=1,3),$
(75) $\tilde{W}_{n,2}(u,\hat{\alpha},\zeta)=\exp\left\{-u\hat{\xi}+\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{\tilde{e}_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\tilde{\eta}_{n,2}(u,\hat{\alpha},\zeta)\right\},$
@@
(76) $\tilde{W}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)=\exp\left\{-u\hat{\xi}+\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{\tilde{e}_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\tilde{\eta}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)\right\}.$
Error bounds and regions of validity are the same as those for the corresponding non-derivative functions, except of course with the different coefficients.
Now, using Eqs. 16 and 50, we can match solutions recessive at $\zeta=\pm\infty$ and $\zeta=\pm i\infty$, and as a result we have
(77) $U^{\prime}\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)=-e^{\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\left\{\frac{1}{8}u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}\tilde{W}_{n,1}(u,\hat{\alpha},\zeta),$
(78) $U^{\prime}\left(\tfrac{1}{2}u\hat{\alpha}^{2},i\sqrt{2u}\,\zeta\right)=-e^{-\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\left\{\frac{1}{8}u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}\tilde{W}_{n,3}(u,\hat{\alpha},\zeta),$
(79) $U^{\prime}\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)=-\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\left\{\frac{1}{8}u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}\tilde{W}_{n,2}(u,\hat{\alpha},\zeta),$
@@
(80) $U^{\prime}\left(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta\right)=\pm ie^{\mp\frac{1}{2}u\hat{\alpha}^{2}\pi i}\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\\ \times\left\{\frac{1}{8}u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}\tilde{W}_{n,4}^{\pm}(u,\hat{\alpha},\zeta).$
@@ 4 Expansions valid at the turning points
(81) $w_{n,1}(u,a,z)=e^{-\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\Biggl{\{}U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)\mathcal{A}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)}{\partial\zeta}\mathcal{B}(u,a,z)\right\},$
@@
(82) $w_{n,3}(u,a,z)=e^{\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\Biggl{\{}U\left(\tfrac{1}{2}u\hat{\alpha}^{2},i\sqrt{2u}\,\zeta\right)\mathcal{A}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(\tfrac{1}{2}u\hat{\alpha}^{2},i\sqrt{2u}\,\zeta\right)}{\partial\zeta}\mathcal{B}(u,a,z)\right\},$
where $w_{n,1}(u,a,z)$ and $w_{n,3}(u,a,z)$ are the solutions defined by Eqs. 18 and 19, which are recessive at the singularities $z=x_{1}$ and $z=x_{3}$, respectively. Note that $U(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta)$ is recessive at $\zeta=i\infty$, which corresponds to $z=x_{1}$, and thus the RHS of (81) matches the recessive behavior of the function on the LHS of the equation. Likewise, both sides of (82) have a matching recessive behavior at $z=x_{3}$ ($\zeta=-i\infty$).
The multiplicative constants on the RHS of Eqs. 81 and 82 were chosen to produce, on referring to Eqs. 45 and 30,
(83) $w_{n,2}(u,a,z)=\frac{\sqrt{2\pi}}{\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)}\Biggl{\{}U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)\mathcal{A}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)}{\partial\zeta}\mathcal{B}(u,a,z)\right\}.$
@@
We need a similar representation for $w_{n,4}(u,a,z)$ that involves $U(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta)$ and its derivative, both of which are recessive at $z=x_{4}$ ($\zeta=-\infty$). This is achieved by defining $\hat{\alpha}$ by
@@
where $\lambda$ is the connection coefficient that is defined via (31). Then from Eqs. 31, 45, 81, 82, and 84 we arrive at our desired representation
(85) $w_{n,4}(u,a,z)=\frac{\sqrt{2\pi}}{\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)}\Biggl{\{}U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta\right)\mathcal{A}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta\right)}{\partial\zeta}\mathcal{B}(u,a,z)\right\}.$
Solving Eqs. 81 and 83, and using the Wronskian (43), yields the explicit formulas
(86) $\mathcal{A}(u,a,z)=\frac{e^{-\frac{1}{4}(u\hat{\alpha}^{2}+1)\pi i}\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)}{2\sqrt{u\pi}}\left\{w_{n,2}(u,a,z)\frac{\partial U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)}{\partial\zeta}\right.\\ \left.+\frac{i}{\sqrt{2u}}w_{n,1}(u,a,z)\frac{\partial U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)}{\partial\zeta}\right\},$
@@
(87) $\mathcal{B}(u,a,z)=-\frac{e^{-\frac{1}{4}(u\hat{\alpha}^{2}+1)\pi i}\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)}{2\sqrt{u\pi}}\left\{w_{n,2}(u,a,z)U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)\right.\\ \left.-\frac{i}{\sqrt{2u}}w_{n,1}(u,a,z)U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)\right\}.$
Similar representations can be obtained by choosing any pair of Eqs. 81, 82, 83, and 85.
The plan is to replace the six functions $w_{n,j}(u,a,z)$ ($j=1,2$), $U(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta)$, $\partial U(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta)/\partial\zeta$, $U(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta)$ and $\partial U(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta)/\partial\zeta$ in Eqs. 86 and 87 by their LG expansions, and likewise in other sectors containing numerically satisfactory sets of functions. This provides a method of asymptotically evaluating the two coefficient functions in certain regions described next, since each product pair in the expression involves one recessive and one dominant solution in the region under consideration.

### v3

@@ 3 LG expansions for parabolic cylinder functions
Here we present relevant results for parabolic cylinder functions, and then derive LG expansions for the ones that we shall use in our uniform asymptotic expansions for solutions of (1). A comprehensive asymptotic study of these functions was given by [6] for the case where the parameter is large. These involved LG and Airy function expansions, which in combination are valid for all values of the complex argument, including those at turning points. Here we consider the case where the parameter of the parabolic cylinder functions, namely ±$\pm\frac{1}{2}u\hat{\alpha}^{2}$, can be both small and large in absolute value. We do not require, or obtain, expansions that are valid at the turning points of the equation they satisfy.
Firstly, the following Wronskians hold (see [3, Eqs. 12.2.11 and 12.2.12]):
@@
Next, connection formulas are given as follows ([3, Eqs. 12.2.13, 12.2.18, 12.2.19 and 12.2.20])
@@
Thus, for the coefficients [4, Eq. (1.9)], let us denote $E_{s}(\xi)=e_{s}(\hat{\alpha},\hat{\beta})$. Then we find, using Eqs. 55, 56, 58, and 59, and [4, Eqs. (1.10) - (1.12)], that
@@
(63) $\hat{W}_{j}(u,\hat{\alpha},\zeta)=\exp\left\{u\hat{\xi}+\sum\limits_{s=1}^{n-1}{\frac{e_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\hat{\eta}_{n,j}(u,\hat{\alpha},\zeta)\right\}\quad(j=1,3),$
(64) $\hat{W}_{2}(u,\hat{\alpha},\zeta)=\exp\left\{-u\hat{\xi}+\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{e_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\hat{\eta}_{n,2}(u,\hat{\alpha},\zeta)\right\},$
@@
(65) $\hat{W}_{4}^{\pm}(u,\hat{\alpha},\zeta)=\exp\left\{-u\hat{\xi}+\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{e_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\hat{\eta}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)\right\}.$
Similarly to (28), the LG solutions $\hat{W}_{4}^{\pm}(u,\hat{\alpha},\zeta)$ only differ by a multiplicative constant, and their $\zeta$ regions of validity $\Theta_{4}^{\pm}$ (where $\hat{\eta}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)=\mathcal{O}(u^{-n})$ as $u\to\infty$) differ.
Note all of these functions are independent of $n$ due to their (unique) behavior at infinity in sectors where they are recessive. For example, since $\hat{\beta}\to 0$ as $|\zeta|\to\infty$, $e_{s}(\hat{\alpha},0)=0$, and from the error bounds provided below $\hat{\eta}_{n,1}(u,i\infty)=0$, we have from Eq. 63
(66) $\hat{W}_{1}(u,\hat{\alpha},\zeta)\sim e^{u\hat{\xi}}\quad(\zeta\to i\infty,\,\Re(\hat{\xi})\to-\infty),$
and similarly for the other LG solutions.
@@
Matching solutions recessive at $\zeta=\pm\infty$ and $\zeta=\pm i\infty$, and using Eqs. 15 and 54, provides the following:
(70) $U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)=\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\frac{e^{\frac{1}{4}(u\hat{\alpha}^{2}+1)\pi i}}{\left\{2u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\hat{W}_{1}(u,\hat{\alpha},\zeta),$
(71) $U\left(\tfrac{1}{2}u\hat{\alpha}^{2},i\sqrt{2u}\,\zeta\right)=\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\frac{e^{-\frac{1}{4}(u\hat{\alpha}^{2}+1)\pi i}}{\left\{2u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\hat{W}_{3}(u,\hat{\alpha},\zeta),$
(72) $U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)=\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\frac{1}{\left\{2u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\hat{W}_{2}(u,\hat{\alpha},\zeta),$
@@
(73) $U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta\right)=\pm ie^{\mp\frac{1}{2}u\hat{\alpha}^{2}\pi i}\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\\ \times\frac{1}{\left\{2u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\hat{W}_{4}^{\pm}(u,\hat{\alpha},\zeta),$
@@
Next, for the derivatives, from (13) (with $\alpha$ replaced by $\hat{\alpha}$), we find that
(74) $\frac{d^{2}\tilde{Y}}{d\zeta^{2}}=\left\{u^{2}\left(\zeta^{2}-\hat{\alpha}^{2}\right)+\frac{2\zeta^{2}+\hat{\alpha}^{2}}{\left(\zeta^{2}-\hat{\alpha}^{2}\right)^{2}}\right\}\tilde{Y},$
has solutions given by
(75) $\tilde{Y}=(\zeta^{2}-\hat{\alpha}^{2})^{-1/2}\partial Y/\partial\zeta,$
where $Y=U(-\frac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta)$, or any other solution of (13) with $\alpha$ replaced by $\hat{\alpha}$. Now, again with $f=\zeta^{2}-\hat{\alpha}^{2}$, but now with $g=(2\zeta^{2}+\hat{\alpha}^{2})/(\zeta^{2}-\hat{\alpha}^{2})^{2}$ instead of $g=0$, we have in place of $\psi(\hat{\alpha},\zeta)$ as given by (55) the Schwarzian, denoted by $\tilde{\psi}(\hat{\alpha},\zeta)$, given by
@@
Then, similarly to Eqs. 60, 61, and 62, we obtain LG coefficients given by
@@
with $\hat{\beta}$ also given by (56), and subsequent coefficients by (62). Again, $\tilde{e}_{2s}(\hat{\alpha},\hat{\beta})$ are meromorphic at $\zeta=\pm\hat{\alpha}$ when regarded as functions of $\zeta$, and $\tilde{e}_{2s+1}(\hat{\alpha},\hat{\beta})$ when regarded as functions of $\zeta$ are analytic at all points away from the interval $[-\hat{\alpha},\hat{\alpha}]$.
The LG solutions to (74) are given by $\tilde{Y}_{j}(u,\hat{\alpha},\zeta)=(\zeta^{2}-\hat{\alpha}^{2})^{-1/4}\tilde{W}_{j}(u,\hat{\alpha},\zeta)$ where
(79) $\tilde{W}_{j}(u,\hat{\alpha},\zeta)=\exp\left\{u\hat{\xi}+\sum\limits_{s=1}^{n-1}{\frac{\tilde{e}_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\tilde{\eta}_{n,j}(u,\hat{\alpha},\zeta)\right\}\quad(j=1,3),$
(80) $\tilde{W}_{2}(u,\hat{\alpha},\zeta)=\exp\left\{-u\hat{\xi}+\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{\tilde{e}_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\tilde{\eta}_{n,2}(u,\hat{\alpha},\zeta)\right\},$
@@
(81) $\tilde{W}_{4}^{\pm}(u,\hat{\alpha},\zeta)=\exp\left\{-u\hat{\xi}+\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{\tilde{e}_{s}(\hat{\alpha},\hat{\beta})}{u^{s}}}\right\}\left\{1+\tilde{\eta}_{n,4}^{\pm}(u,\hat{\alpha},\zeta)\right\},$
again all being independent of $n$. Error bounds and regions of validity are the same as those for the corresponding non-derivative functions, except of course with the different coefficients.
Now, using Eqs. 16 and 54, we can match solutions that are recessive at $\zeta=\pm\infty$ and $\zeta=\pm i\infty$, and as a result we arrive at our desired expansions:
(82) $U^{\prime}\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)=-e^{\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\left\{\frac{1}{8}u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}\tilde{W}_{1}(u,\hat{\alpha},\zeta),$
(83) $U^{\prime}\left(\tfrac{1}{2}u\hat{\alpha}^{2},i\sqrt{2u}\,\zeta\right)=-e^{-\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\left\{\frac{1}{8}u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}\tilde{W}_{3}(u,\hat{\alpha},\zeta),$
(84) $U^{\prime}\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)=-\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\left\{\frac{1}{8}u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}\tilde{W}_{2}(u,\hat{\alpha},\zeta),$
@@
(85) $U^{\prime}\left(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta\right)=\pm ie^{\mp\frac{1}{2}u\hat{\alpha}^{2}\pi i}\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\\ \times\left\{\frac{1}{8}u\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}\tilde{W}_{4}^{\pm}(u,\hat{\alpha},\zeta).$
@@ 4 Expansions valid at the turning points
(86) $w_{1}(u,a,z)=e^{-\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\Biggl{\{}U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)\mathcal{A}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)}{\partial\zeta}\mathcal{B}(u,a,z)\right\},$
@@
(87) $w_{3}(u,a,z)=e^{\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\Biggl{\{}U\left(\tfrac{1}{2}u\hat{\alpha}^{2},i\sqrt{2u}\,\zeta\right)\mathcal{A}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(\tfrac{1}{2}u\hat{\alpha}^{2},i\sqrt{2u}\,\zeta\right)}{\partial\zeta}\mathcal{B}(u,a,z)\right\},$
where $w_{1}(u,a,z)$ and $w_{3}(u,a,z)$ are the solutions defined by Eqs. 18 and 19, which are recessive at the singularities $z=x_{1}$ and $z=x_{3}$, respectively. We note that $U(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta)$ is recessive at $\zeta=i\infty$, which corresponds to $z=x_{1}$, and thus the RHS of (86) matches the recessive behavior of the function on the LHS of the equation. Likewise, both sides of (87) have a matching recessive behavior at $z=x_{3}$ ($\zeta=-i\infty$).
The multiplicative constants on the RHS of Eqs. 86 and 87 were chosen to produce, on referring to Eqs. 49 and 30,
(88) $w_{2}(u,a,z)=\frac{\sqrt{2\pi}}{\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)}\Biggl{\{}U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)\mathcal{A}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)}{\partial\zeta}\mathcal{B}(u,a,z)\right\}.$
@@
We need a similar representation for $w_{4}(u,a,z)$ that involves $U(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta)$ and its derivative, both of which are recessive at $z=x_{4}$ ($\zeta=-\infty$). This is achieved by defining $\hat{\alpha}$ by
@@
where $\lambda$ is the connection coefficient that is defined via (31). Then from Eqs. 31, 49, 86, 87, and 89 we arrive at our desired representation
(90) $w_{4}(u,a,z)=\frac{\sqrt{2\pi}}{\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)}\Biggl{\{}U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta\right)\mathcal{A}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},-\sqrt{2u}\,\zeta\right)}{\partial\zeta}\mathcal{B}(u,a,z)\right\}.$
Solving Eqs. 86 and 88, and using the Wronskian (47), yields the explicit formulas
(91) $\mathcal{A}(u,a,z)=\frac{e^{-\frac{1}{4}(u\hat{\alpha}^{2}+1)\pi i}\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)}{2\sqrt{u\pi}}w_{2}(u,a,z)\frac{\partial U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)}{\partial\zeta}\\ +\frac{i}{\sqrt{2u}}w_{1}(u,a,z)\frac{\partial U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right)}{\partial\zeta},$
@@
(92) $\mathcal{B}(u,a,z)=-\frac{e^{-\frac{1}{4}(u\hat{\alpha}^{2}+1)\pi i}\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)}{2\sqrt{u\pi}}w_{2}(u,a,z)U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)\\ -\frac{i}{\sqrt{2u}}w_{1}(u,a,z)U\left(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta\right).$
Similar representations can be obtained by choosing any pair of Eqs. 86, 87, 88, and 90.
The plan is to replace the six functions $w_{j}(u,a,z)$ ($j=1,2$), $U(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta)$, $\partial U(-\tfrac{1}{2}u\hat{\alpha}^{2},\sqrt{2u}\,\zeta)/\partial\zeta$, $U(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta)$ and $\partial U(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta)/\partial\zeta$ in Eqs. 91 and 92 by their LG expansions, and likewise in other sectors containing numerically satisfactory sets of functions. This provides a method of asymptotically evaluating the two coefficient functions in certain regions described next, since each product pair in the expression involves one recessive and one dominant solution in the region under consideration.